OpenStudy (anonymous):

When do you use sine, cosine, and tangent, and when do you use their reciprocals of these also known as cosine, secant, and cotangent

OpenStudy (anonymous):

Anyone please

OpenStudy (zzr0ck3r):

this is like asking when do you use 2, and 3, and when do you use 1/3 and 1/2

OpenStudy (anonymous):

Well can you explain

OpenStudy (anonymous):

I'm confused

OpenStudy (zzr0ck3r):

there are 10000000000000.............. different ways to use them

OpenStudy (zzr0ck3r):

can you be more specific?

OpenStudy (anonymous):

I know using a negative sine is when you are looking for an angle

OpenStudy (anonymous):

I'm confused on this question whether to use cosine or secant

OpenStudy (zzr0ck3r):

um, I think you may need to read about the unit circle and what sine and cosine are, negative sine does not imply you are looking for an angle.

OpenStudy (zzr0ck3r):

what is the question?

OpenStudy (anonymous):

The picture below shows a right triangle shaped charging stand for a gaming system. Which expression shows the length of side AB? 8 cosec 60° 8/cos60°

OpenStudy (anonymous):

I know it's either one of those but am unsure

OpenStudy (zzr0ck3r):

soh cah toa cos(x) = adjacent over hypotenuse so let so cos(60) = 8/AB so AB = 8/cos(60)

OpenStudy (anonymous):

How is it not cosecant that's where I'm confused

OpenStudy (anonymous):

ANYONE?

OpenStudy (zzr0ck3r):

I don't know why it would be again soh cah toa cah implies cos = adjacent/hypot0nuse I have not said the word cosecant.....

OpenStudy (zzr0ck3r):

a = b/c you are given a and b solve for c

OpenStudy (anonymous):

I know but my original question was knowing when to use secant

OpenStudy (zzr0ck3r):

when you don't want to write 1/cos

OpenStudy (anonymous):

So I don't know which one to use

OpenStudy (anonymous):

How would I know not to make it in hypotenuse over adjacent

OpenStudy (zzr0ck3r):

1/cos(60) = sec(6) 1/sin(60) = csc(60)

OpenStudy (zzr0ck3r):

you remember the rule

OpenStudy (zzr0ck3r):

soh cah toa

OpenStudy (anonymous):

What does it mean

OpenStudy (zzr0ck3r):

 soh means sin(theta) = opposite / hypotenuse cah means cos(theta) = adjacent/hypotenuse toa means tan(theta) = opposite/adjacent

OpenStudy (anonymous):

You are avoiding my original question

OpenStudy (zzr0ck3r):

lol

OpenStudy (zzr0ck3r):

ok when do you use the number 2?

OpenStudy (anonymous):

What?

OpenStudy (zzr0ck3r):

exactly

OpenStudy (anonymous):

WHEN DO YOU USE SECANT!!!!!!!!!!!!!!

OpenStudy (zzr0ck3r):

WHEN DO YOU USE 2!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

AND ALL OF THE OTHER RECIPROCALS

OpenStudy (zzr0ck3r):

secant(x) = 1/cos(x)

OpenStudy (zzr0ck3r):

that's all I can tell you

OpenStudy (anonymous):

I know how to do it

OpenStudy (anonymous):

Just want to know when

OpenStudy (anonymous):

So in the future I will not need help

OpenStudy (zzr0ck3r):

ok say you have something like cos(x) = 1/2 then 2 = 1/cos(x) maybe you don't like to write fractions so you write it as 2 = sec(x)

OpenStudy (anonymous):

When you are solving an equation like the one I showed you above

OpenStudy (zzr0ck3r):

if you don't understand that I don't know what to tell you.....and using all caps will not help. You are asking a VERY open ended question. Your question is similar to me asking you when do you use 2 one can use 2 in many situations...

OpenStudy (zzr0ck3r):

again when ever you have 1/cos(x) you use it

OpenStudy (zzr0ck3r):

so again

OpenStudy (zzr0ck3r):

if you have 1/cos(x) write it as sec(x)

OpenStudy (anonymous):

Never mind you were no help

OpenStudy (anonymous):

But I will medal you for trying

OpenStudy (zzr0ck3r):

listen you need to grow up, and learn to ask a question. Don't be ruse. I am taking my time to sit hear and try and explain something to you....

OpenStudy (zzr0ck3r):

I don't need a medal

OpenStudy (zzr0ck3r):

there 5billion different ways you would use secant in a trig question. Your question is to vague..............

OpenStudy (anonymous):

Okay zzr0ck I learned that we solve using the reciprocal as well when you are solving a problem

OpenStudy (anonymous):

Besides the course I am taking is Geometry

OpenStudy (anonymous):

So it will be very basic trigonometry

OpenStudy (zzr0ck3r):

one example  cos(x) = 2/4 so 4 = 2/cos(x) = 2sec(x)

OpenStudy (anonymous):

Go back to my question then why is the other answer incorrect

OpenStudy (zzr0ck3r):

because the answer is 8/cos(60) the other answer says 8csc(60) = 8/sin(60)

OpenStudy (anonymous):

OOOOHHHHH

OpenStudy (anonymous):

THANK YOU VERY MUCH

OpenStudy (zzr0ck3r):

so the answer is 8/cos(60) you may also write this as 8sec(60)

OpenStudy (anonymous):

Didn't see that

OpenStudy (anonymous):

I'm very sorry for the trouble that I caused

OpenStudy (anonymous):

Along with my fit

OpenStudy (zzr0ck3r):

so your question, after all this time was does 8/cos(60) = 8cosecant(60)

OpenStudy (anonymous):

Haha I guess so